# The Roots of Performance Disparity in Multilingual Language Models: Intrinsic Modeling Difficulty or Design Choices?

### v1

@@ 2.2 Morphological Complexity
However, later analyses show that these correlations often reflect modeling artifacts rather than inherent difficulty. Large-scale multilingual studies demonstrate that apparent morphology effects are confounded with tokenization quality, dataset size, and encoding inefficiencies (Mielke et al., 2019). Morphology-aware segmentation (e.g., Morfessor) substantially reduces surprisal gaps induced by standard BPE (Park et al., 2021; Mager, 2022; Saleva and Lignos, 2021), and analyses of WordPiece/BPE segmentations show inadequacies for complex morphology and derivation (Klein and Tsarfaty, 2020; Lerner and Yvon, 2025).
@@ 3.1 Tokenization: From Frequency-Based Segments to Linguistically Informed Units
Tokenization is a widely studied design lever. Frequency-driven subword algorithms (BPE, WordPiece) often fragment morphology and non-Latin scripts, inflating sequence length, training costs, and obscuring linguistically meaningful units (Park et al., 2021; Mielke et al., 2019; Ali et al., 2024a; Lundin et al., 2025; Gutherz et al., 2023; Ali et al., 2024b). Across studies, segmentation quality explains a substantial portion of cross-linguistic gaps, where morphology- and script-aware methods reduce surprisal in agglutinative and polysynthetic languages (Creutz and Lagus, 2007; Ling et al., 2015; Mager, 2022; Saleva and Lignos, 2021; Loukatou et al., 2021; Dang et al., 2025).
@@
Implication 1: Use byte-/character-level models, and treat script-specific tokenizers as first-class components (Qarah and Alsanoosy, 2024; Alrefaie et al., 2024; Rana et al., 2025). Adaptive and multilingual approaches that balance scripts and allocate shared-vocabulary capacity can reduce parity gaps without compromising high-resource performance (Bostrom and Durrett, 2020; Limisiewicz et al., 2023; Ahia et al., 2024; Baxi and Bhatt, 2024). Byte-level and tokenizer-free models show competitive performance at scale, trading longer sequences for reduced script penalties (Kenter et al., 2018; Pagnoni et al., 2025). We note that byte-level models are more expensive, and thus might be less usable for a fixed budget.

### v2

@@ 2.2 Morphological Complexity
However, later analyses show that these correlations often reflect modeling artifacts rather than inherent difficulty. Large-scale multilingual studies demonstrate that apparent morphology effects are confounded with tokenization quality, dataset size, and encoding inefficiencies (Mielke et al., 2019). Morphology-aware segmentation (e.g., Morfessor) substantially reduces surprisal gaps induced by standard BPE (Park et al., 2021; Mager et al., 2022; Saleva and Lignos, 2021), and analyses of WordPiece/BPE segmentations show inadequacies for complex morphology and derivation (Klein and Tsarfaty, 2020; Lerner and Yvon, 2025).
@@ 3.1 Tokenization: From Frequency-Based Segments to Linguistically Informed Units
Tokenization is a widely studied design lever. Frequency-driven subword algorithms (BPE, WordPiece) often fragment morphology and non-Latin scripts, inflating sequence length, training costs, and obscuring linguistically meaningful units (Park et al., 2021; Mielke et al., 2019; Ali et al., 2024a; Lundin et al., 2025; Gutherz et al., 2023; Ali et al., 2024b). Across studies, segmentation quality explains a substantial portion of cross-linguistic gaps, where morphology- and script-aware methods reduce surprisal in agglutinative and polysynthetic languages (Creutz and Lagus, 2007; Ling et al., 2015; Mager et al., 2022; Saleva and Lignos, 2021; Loukatou et al., 2021; Dang et al., 2025).
@@
Implication 1: Use byte-/character-level models, and treat script-specific tokenizers as first-class components (Qarah and Alsanoosy, 2024; Alrefaie et al., 2024; Rana et al., 2025). Adaptive and multilingual approaches that balance scripts and allocate shared-vocabulary capacity can reduce parity gaps without compromising high-resource performance (Bostrom and Durrett, 2020; Limisiewicz et al., 2023; Ahia et al., 2024; Baxi and Bhatt, 2024). Byte-level and tokenizer-free models show competitive performance at scale, trading longer sequences for reduced script penalties (Pagnoni et al., 2025). We note that byte-level models are more expensive, and thus might be less usable for a fixed budget.